The Technology Company issues $506,000 of 10%, 10−year bonds at 108 on March 31, 2018. The bonds pay interest on March 31 and September 30. Assume that the company uses the straight−line method for amortization. Calculate the net balance that will be reported for the bonds on the September 30, 2018 balance sheet. (Round your intermediate answers to the nearest dollar.) a. $506,000 b. $546,480 c. $544,456 d. $548, 504
Answer #1

Answer : C) $ 544,456 (Answer)

Face value of the bond = $ 506,000

Issued at @108 = 506,000*1.08 = $ 546,480

Premium on Bonds Payable = $ 40,480

Number of parments = 10*2 = 20

semi- annual amortization of bonds payable = $40,480 / 20 = 2,024

carrying valueon bonds Payable : $ 546,480- 2024 = $ 544,456 ( Answer)